OpenStudy (anonymous):

Medal A scalpel blade is to be 3/8 in long. How many blades can be cut from a piece of stainless steel that is 3 1/2 ft long? ? blades

jimthompson5910 (jim_thompson5910):

3/8 inches = 0.375 inches 3 ft + 1/2 ft = 3 ft + 0.5 ft = 3.5 ft 3.5 ft = (3.5 ft)*(12 inches/1 ft) = 42 inches

jimthompson5910 (jim_thompson5910):

so the question is asking "A scalpel blade is to be 0.375 in long. How many blades can be cut from a piece of stainless steel that is 42 inches long? "

OpenStudy (anonymous):

15.75

OpenStudy (anonymous):

i think you are right

jimthompson5910 (jim_thompson5910):

it asks "How many blades", so it's assumed that you have a whole number of blades

OpenStudy (anonymous):

so 16

jimthompson5910 (jim_thompson5910):

x = # of blades (# of blades)*(length of each blade) = total length of steel x*0.375 = 42 0.375x = 42 solve for x

OpenStudy (anonymous):

42

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

42-0.375 = 41.625

jimthompson5910 (jim_thompson5910):

you don't subtract

jimthompson5910 (jim_thompson5910):

0.375x means 0.375 times x you need to undo that multiplication to isolate x

OpenStudy (anonymous):

0.140

jimthompson5910 (jim_thompson5910):

idk how you got that

OpenStudy (anonymous):

0.375*0.375

jimthompson5910 (jim_thompson5910):

I see. That is incorrect.

OpenStudy (anonymous):

so what do i need to mulitply

jimthompson5910 (jim_thompson5910):